OpenStudy (anonymous):

Aritmetic QS. if a,b,c,d are real numbers under what condition we have [a,b] are subset of (cd)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

maybe,,,,, a and b are subset of c and d only if the values of a are not smaller that c and the values of b are not bigger than d.

hero (hero):

If a + b = c a - b = d Just taking a stab at it....

OpenStudy (anonymous):

lo,, but why u perform operations?

hero (hero):

To relate a and b to c and d

hero (hero):

Now you can describe a and b in terms of c and d. Now that we've done that.... a and b are components of c and d

OpenStudy (anonymous):

mmm.. so the condition where a and be are subset of c and d is when performing an operation between a and b our result will be either c or d.

OpenStudy (amistre64):

does that read; on the interval between a and b, under what conditions do we get a product of c and d?

OpenStudy (anonymous):

nop

OpenStudy (amistre64):

i cant make sense of the notation :)

OpenStudy (anonymous):

under what what conditions we have [ab] as a subset of (c,d)

hero (hero):

Is there a graph associated with this? If so, use the drawing pad to create it.

OpenStudy (anonymous):

\[[a,b] \subseteq (c,d)\]

OpenStudy (amistre64):

it appears to be interval notation

OpenStudy (anonymous):

no,,the professor just gave it like that yes they are interval

OpenStudy (anonymous):

:(

OpenStudy (amistre64):



OpenStudy (anonymous):

yesss kind offffff.. is what i have in my mind

OpenStudy (amistre64):

when the |a+b| < |c+d| perhaps?

OpenStudy (amistre64):

say we got (1 [3 4] 2) ; 7 > 3 tho

OpenStudy (anonymous):

yess so i sayd that a and b are subset of c and d only if the values of a are not smaller that c and the values of b are not bigger than d.

OpenStudy (amistre64):

(1 [2 3 4 5 6 7 ] 8); 8+1 = 9, 7+2 = 9 6+3 = 9 7+6 = 13 2+3 = 5 hmmm

OpenStudy (amistre64):

needs some work to it, but i think its along those lines yes

OpenStudy (amistre64):

|8-1| = 7 |2-7| = 5

OpenStudy (amistre64):

that seems to work better; if the absolute value between a and b is smaller than c to d; then it is a subset? getting closer

OpenStudy (anonymous):

ohh there u include the absule value

OpenStudy (amistre64):

yes, i would since its the distance between points that is important and not direction

OpenStudy (amistre64):

If |a-b| < |c-d| then [a,b] is a proper subset of (c,d)

OpenStudy (anonymous):

ok..

OpenStudy (amistre64):

might have to check the boundaries tho to make sure;

OpenStudy (amistre64):

spose you have (3,8) and [2,4]; that fits what I wrote, but doesnt work

OpenStudy (amistre64):

if c<d and a<d and b<d; and a>c and b>c would be restrictions

OpenStudy (anonymous):

noooo because i think 3 and 8 must be elements between z and 4

OpenStudy (anonymous):

so should be [3,8] and (2,10)

OpenStudy (anonymous):

if i go with ur definition

OpenStudy (amistre64):

if you go with my definition; you have to specify the relationship between a b c and d so that you arent taking values that are "off limits"

OpenStudy (anonymous):

no he said they r real numbers... so they are infinitely whatever value

OpenStudy (amistre64):

c <a < b < d

OpenStudy (amistre64):

they can still be real numbers; but under certain conditions; otherwise there is no way to say that [a,b] is a subset of (c,d) if we can any combonation of numbers

OpenStudy (anonymous):

:(... so i need to find out the boundaries???

OpenStudy (amistre64):

[4,19] is not a subset of (-5,3)

OpenStudy (amistre64):

you need to define boundaries yes

OpenStudy (amistre64):

not so much boundaries, more of conditions

OpenStudy (amistre64):

There exists some form of interval that can be defined in this manner with these restrictions such that one is a subset of the other

OpenStudy (anonymous):

where do u think i can read more about it

OpenStudy (amistre64):

whats the topic?

OpenStudy (anonymous):

I'm not sure I'm following this, but, as I understand it, [a,b] is a subset of (c,d) if c<a<d and c<b<d. Shouldn't that be enough?

OpenStudy (amistre64):

set theory perhaps, but i aint got no dedicated sources for it

OpenStudy (amistre64):

it might be, but then again it might not be :)